OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations auditioned today? 430 720 1150 1300

OpenStudy (shane_b):

Let x be the number of one-step equations and y be the number of two-step equations: x+y=1150 3x-2y=1300 Can you solve it from here?

OpenStudy (anonymous):

no i dont get it.

OpenStudy (shane_b):

Do you at least understand how the equations above were generated? If so, I'll show you the rest.

OpenStudy (anonymous):

you put it all together from the word problem

OpenStudy (shane_b):

Yes, if you understand how I did that, we can move on to solve the system of equations. That part is pretty easy in this problem.

OpenStudy (shane_b):

Ok...moving on. We developed two equations which have two unknowns: x+y=1150 3x-2y=1300 Our goal is to solve for x and y. I believe the simplest way to do that is to solve one of the equations for a single variable and then substitute that back into the other equation. Looking at the first equation, let's rearrange the terms: \[x+y=1150 ==>x=1150-y\] Now take the second equation and plug in "1150-y" anywhere you see an x: \[3x-2y=1300==>3(1150-y)-2y=1300\]With me so far?

OpenStudy (shane_b):

Guess not :) If I'm here when you get back we can finish this...or someone else can.

OpenStudy (anonymous):

i would write the same answer with @Shane_B